OpenStudy (anonymous):

solve 48 +72 divided by 9

Parth (parthkohli):

Is it \(48 + {72 \over 9}\) or what?

OpenStudy (anonymous):

yes. do i divide 9 into 72 first, then add 48 to the answer?

Parth (parthkohli):

Have you heard of BODMAS? According to that, you divide first.

OpenStudy (lgbasallote):

i am assuming B- bracket..but what's o?

Parth (parthkohli):

So, it'd be 48 + 8 => 56

OpenStudy (anonymous):

no. please update me on BODMAS

Parth (parthkohli):

BODMAS => Brackets of Division, Multiplication, Addition and Subtraction. If brackets aren't given, then you've got to put them in the order given.

OpenStudy (lgbasallote):

another acronym you may wanna know is PEMDAS kinda like it... Parenthesis Exponent Multiplication Division Addition Subtraction

OpenStudy (anonymous):

56.

OpenStudy (anonymous):

no brackets so division comes before addition. now, in the problem 5 x 2 + 8, I would multiply 5 x2 first then add 8, right?

Parth (parthkohli):

But the problem stated above is different. But yeah, you first do division, then multiplication, then addition, then subtraction.

OpenStudy (anonymous):

its nt right

OpenStudy (anonymous):

y u hav mutiple 5 with 2

Parth (parthkohli):

@lindamilesj Can you repeat the problem please?

OpenStudy (anonymous):

48 + (72 divided by 9) = 56

Parth (parthkohli):

Exactly, so it's be (5 * 2) + 8 = 10 + 8 = 56

Parth (parthkohli):

@shruti He's asking about a different problem

OpenStudy (anonymous):

i'am not getting his query.

Parth (parthkohli):

He's telling to solve \(5 * 2 + 8\)

OpenStudy (anonymous):

then it should be 18.

Parth (parthkohli):

Exactly what I stated :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hey guys...first of all I am a lady and secondly, I need the steps, not just the answer so I did the problem like this...5 x 2, first and the sum is 10, then add 8 to it, right?

OpenStudy (anonymous):

Next problem 6 x 3 / 2 + 13

Parth (parthkohli):

\(\Large \color{purple}{ [6(3/2)]+13 }\) \(\Large \color{purple}{ [6(1.5)] + 13 }\) \(\Large \color{purple}{ 9 + 13 }\) \(\Large \color{purple}{ 24 }\)

Parth (parthkohli):

Hey, if you think I helped you, a medal would help me. Thanks :)

Parth (parthkohli):

3/2 = 1.5 Think about it.

Parth (parthkohli):

First, I divided 3 with 2 which resulted in 1.5

OpenStudy (anonymous):

oh, okay. now, here's another one...10/5 + 20/10=

Parth (parthkohli):

\(\Large \color{purple}{ (10 \div 5) + (20 \div 10) }\) \(\Large \color{purple}{ 2 + 2 }\) \(\Large \color{purple}{ 4 }\)

OpenStudy (anonymous):

hey 9 + 13 = 22

Parth (parthkohli):

Oh lol sorry

OpenStudy (anonymous):

okay...25/5-5

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow (25 \div 5) - 5 }\) \(\Large \color{purple}{\rightarrow 5 - 5 }\) \(\Large \color{purple}{\rightarrow 0 }\)

OpenStudy (anonymous):

I got zero for this one because I divided 5 into 25 then subtracted 5 from 5

Parth (parthkohli):

Yes :)

OpenStudy (anonymous):

here's a real one..-4 +(-6) /(2 x 3)

OpenStudy (anonymous):

okay this is the answer I have -4 + (-6)(+6)=-36 + -4 = -40

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow -4 + (-6) \div 6 }\) \(\Large \color{purple}{\rightarrow -4 - 1 }\) \(\Large \color{purple}{\rightarrow -5 }\) If there are brackets, then solve them first.

OpenStudy (anonymous):

no, the problem is (-4)-(-6) x 3/2= so, I think that -4-(-6) x 1.5

OpenStudy (anonymous):

I am basing my answer on the answer you gave before when you divide first so I divided 3 by 2 first...